# Khoroshkin–Tolstoy approach for quantum superalgebras

### v2

@@ 1. Introduction
We start with the necessary information on the nontwisted affine Lie superalgebra ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$. More information one can find in the papers [50, 51].
We denote by $\mathcal{L}(\mathfrak{sl}_{M|N})$ the loop superalgebra of $\mathfrak{sl}_{M|N}$, by ~$\widetilde{\mathcal{L}}(\mathfrak{sl}_{M|N})$ its standard central extension by a one-dimensional center $\mathbb{C}c$, and by ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$ the Lie superalgebra obtained from ~$\widetilde{\mathcal{L}}(\mathfrak{sl}_{M|N})$ by adding a natural derivation $d$. By definition
@@
and we use as the Cartan subsuperalgebra of ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$ the space [^5] Here we use the natural embedding of $\mathfrak{sl}_{M|N}$ into ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$.
@@
We assume that $[c]=[d]=\overline{0}$.
It is convenient to denote
$\widetilde{\mathfrak{h}}=\mathfrak{h}\oplus\mathbb{C}\,c$
and identify $\mathfrak{h}$ and $\widetilde{\mathfrak{h}}$ with the corresponding subspaces of $\widehat{\mathfrak{h}}$. We identify the space $\mathfrak{h}^{*}$ with the subspace of $\widetilde{\mathfrak{h}}^{*}$ defined as
$\{\lambda\in\widetilde{\mathfrak{h}}^{*}\mid\langle\lambda,\,c\rangle=0\}.$
and with the subspace of $\widehat{\mathfrak{h}}^{*}$ defined as
$\{\lambda\in\widehat{\mathfrak{h}}^{*}\mid\langle\lambda,\,c\rangle=0,\ \langle\lambda,\,d\rangle=0\}.$
We define
@@
where we define $d_{0}=1$.
After all we denote by $\delta$ the element of $\widehat{\mathfrak{h}}^{*}$ defined by the equations
@@
The equation
$A_{ij}=\langle\alpha_{j},\,h_{i}\rangle,\qquad i,j=0,1,\ldots,\,M+N-1,$
[…]
$D^{(1)}=\mathrm{diag}(d_{0},d_{1},\ldots,d_{M},d_{M+1},\ldots,d_{M+N-1})=\mathrm{diag}(\underbracket{1,\,\ldots,1}_{M+1},\,\underbracket{-1,\,\ldots,\,-1}_{N-1}),$ (4.1)
we see that the matrix $B^{(1)}=D^{(1)}A^{(1)}$ is symmetric. We take $\alpha_{i}$, $i=0,1,\ldots,M+N-1$, as the simple roots. The full system of positive roots of the Lie superalgebra ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$ is
@@
We fix a non-degenerate symmetric bilinear form on $\widehat{\mathfrak{h}}$ by the equations
@@
Let again $\hbar$ be a nonzero complex number such that $q=\exp\hbar$ is not a root of unity. The quantum loop superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is a unital associative $\mathbb{C}$-algebra generated by the elements
$q^{x},\quad x\in\widetilde{\mathfrak{h}},\qquad e_{i},\quad f_{i},\quad i=0,1,\ldots,M+N-1,.$
The $\mathbb{Z}_{2}$-grading of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is defined on generators as
@@
is called the root lattice of $\mathfrak{gl}_{M|N}$. We denote by $\widehat{Q}_{+}$ and $\widehat{Q}_{-}$ the sublattices
$\widehat{Q}_{+}=\bigoplus_{0=1}^{M+N-1}\mathbb{Z}_{\geq 0}\,\alpha_{i},\qquad\widehat{Q}_{-}=\bigoplus_{i=0}^{M+N-1}\mathbb{Z}_{\leq 0}\,\alpha_{i}.$
Assuming that
@@
we endow $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ with a $\widehat{Q}$-grading. We define the $q$-supercommutator for the quantum loop superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ in the same way as it was done for the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ in section 3.
The defining relations of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ look as [20]
@@
$[\![e_{i},\,e_{j}]\!]=0,\qquad[\![f_{i},\,f_{j}]\!]=0,\qquad\mbox{ if }B_{ij}=0,$ (4.5)
$[\![e_{i},\,[\![e_{i},\,e_{i\pm 1}]\!]]\!]=0,\qquad[\![f_{i},\,[\![f_{i},\,f_{i\pm 1}]\!]]\!]=0,\qquad\mbox{ if }B_{ii}\neq 0,$ (4.6)
$[\![[\![[\![e_{M-1},\,e_{M}]\!],\,e_{M+1}]\!],\,e_{M}]\!]=0,\qquad[\![[\![[\![e_{1},\,e_{0}]\!],\,e_{M+N-1}]\!],\,e_{0}]\!]=0,$ (4.7)
$[\![[\![[\![f_{M-1},\,f_{M}]\!],\,f_{M+1}]\!],\,f_{M}]\!]=0,\qquad[\![[\![[\![f_{1},\,f_{0}]\!],\,f_{M+N-1}]\!],\,f_{0}]\!]=0..$ (4.8)
In the case of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{2|1}))$ we have additional relations
@@
$[\![f_{0},\,[\![f_{2},\,[\![f_{0},\,[\![f_{2},\,f_{1}]\!]]\!]]\!]]\!]=[\![f_{2},\,[\![f_{0},\,[\![f_{2},\,[\![f_{0},\,f_{1}]\!]]\!]]\!]]\!].$
From the point of view of quantum integrable systems, it is important that the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is a Hopf algebra with respect to appropriately defined co-multiplication $\Delta$, antipode $S$ and counit $\varepsilon$. Note that the definition of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is invariant under the replacement of $q$ by $q^{-1}$, in contrast to the definition of the Hopf algebra structure. To eliminate the arising ambiguity, we give the explicit form of the Hopf algebra operations adopted in this paper,
@@ 5. Cartan–Weyl generators
Having in mind the $\widehat{Q}$-grading of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$, we say that an element $a$ of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is a root vector corresponding to a root $\gamma\in\widehat{\Delta}$ if $a\in\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))_{\gamma}$. In particular, the generators $e_{i}$ and $f_{i}$ are root vectors corresponding to the roots $\alpha_{i}$ and $-\alpha_{i}$.
One can construct linearly independent root vectors corresponding to all roots from $\widehat{\Delta}\subset\widehat{Q}$, see, for example, the papers [24, 25, 26]. They are called the Cartan–Weyl generators of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. If some ordering of roots is chosen, then appropriately ordered monomials constructed from these vectors form a Poincaré–Birkhoff–Witt basis of $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. To construct root vectors we follow the procedure, based on normal ordering of positive roots, proposed by Khoroshkin and Tolstoy [24, 25, 26].
@@
Assume that some normal ordering of positive roots is chosen. We say that a pair $(\alpha,\,\beta)$ of positive roots generates a root $\gamma$ if $\gamma=\alpha+\beta$ and $\alpha\prec\beta$. A pair of positive roots $(\alpha,\,\beta)$ generating a root $\gamma$ is said to be minimal if there is no other pair of positive roots $(\alpha^{\prime},\,\beta^{\prime})$ generating $\gamma$ such that $\alpha\prec\alpha^{\prime}\prec\beta^{\prime}\prec\alpha$.
It is convenient to denote a root vector corresponding to a positive root $\gamma$ by $e_{\gamma}$, and a root vector corresponding to a negative root $-\gamma$ by $f_{\gamma}$. Following [24, 25, 26], we define root vectors by the following inductive procedure. Given a root $\gamma\in\widehat{\Delta}_{+}$, let $(\alpha,\,\beta)$ be a minimal pair of positive roots generating $\gamma$. Now, if the root vectors $e_{\alpha}$, $e_{\beta}$ and $f_{\alpha}$, $f_{\beta}$ are already constructed, we define the root vectors $e_{\gamma}$ and $f_{\gamma}$ as
$e_{\gamma}\propto[\![e_{\alpha}\,,\,e_{\beta}]\!],\qquad f_{\gamma}\propto[\![f_{\alpha}\,,\,f_{\beta}]\!].$
@@
We use the normal order of the positive roots of ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$ defined as follows. First order the positive roots of $\mathfrak{sl}_{M|N}$ assuming that $\alpha_{ij}\prec\alpha_{kl}$ if $i<k$, or if $i=k$ and $j<l$. Then, $\alpha+m\delta\prec\beta+n\delta$, with $\alpha,\beta\in\Delta_{+}$ and $m,n\in\mathbb{Z}_{\geq 0}$, if $\alpha\prec\beta$, or $\alpha=\beta$ and $m<n$. Further, $(\delta-\alpha)+m\delta\prec(\delta-\beta)+n\delta$, with $\alpha,\beta\in\Delta_{+}$, if $\alpha\prec\beta$, or $\alpha=\beta$ and $m>n$. Finally, we assume that the relation (5.1) is valid.
@@
$e_{\alpha_{ij}}=[\![e_{\alpha_{i,\,j-1}},\,e_{\alpha_{j-1,\,j}}]\!],\qquad f_{\alpha_{ij}}=[\![f_{\alpha_{i,\,j-1}},\,f_{\alpha_{j-1,\,j}}]\!]$ (5.2)
for $j-i>1$. Further, taking into account that $\alpha_{0}=\delta-\alpha_{1,\,M+N}$, we put
@@
$e_{\delta-\alpha_{i,\,M+N}}=[\![e_{\alpha_{i-1,\,i}},\,e_{\delta-\alpha_{i-1,\,M+N}}]\!],\qquad f_{\delta-\alpha_{i,\,M+N}}=[\![f_{\alpha_{i-1,\,i}},\,f_{\delta-\alpha_{i-1,\,M+N}}]\!]$ (5.4)
@@
$e_{\delta-\alpha_{ij}}=[\![e_{\alpha_{j,\,j+1}},\,e_{\delta-\alpha_{i,\,j+1}}]\!],\qquad f_{\delta-\alpha_{ij}}=[\![f_{\alpha_{j,\,j+1}},\,f_{\delta-\alpha_{i,\,j+1}}]\!]$ (5.5)
for $j<M+N$. The root vectors corresponding to the roots $\delta$ and $-\delta$ are indexed by the simple roots $\alpha_{i}$, and are defined by the relations
$e^{\prime}_{\delta;\,\alpha_{i}}=(-1)^{[\alpha_{i}]}[\![e_{\alpha_{i}},\,e_{\delta-\alpha_{i}}]\!],\qquad f^{\prime}_{\delta;\,\alpha_{i}}=(-1)^{[\alpha_{i}]}[\![f_{\alpha_{i}},\,f_{\delta-\alpha_{i}}]\!].$ (5.6)
@@
$f_{\alpha_{ij}+n\delta}=\frac{(-1)^{[\alpha_{i}]}}{[(\,\alpha_{ij}\,|\,\alpha_{i}\,)]_{q}}[\![f_{\alpha_{ij}+(n-1)\delta},\,f^{\prime}_{\delta;\,\alpha_{i}}]\!],$ (5.9)
@@
$f_{(\delta-\alpha_{ij})+n\delta}=\frac{(-1)^{[\alpha_{i}]}}{[(\,\alpha_{ij}\,|\,\alpha_{i}\,)]_{q}}[\![f^{\prime}_{\delta;\,\alpha_{i}},\,f_{(\delta-\alpha_{ij})+(n-1)\delta}]\!],$ (5.11)
@@
$f_{\alpha_{ij}+n\delta}=\frac{(-1)^{[\alpha_{i-1}]}}{[(\,\alpha_{ij}\,|\,\alpha_{i-1}\,)]_{q}}[\![f_{\alpha_{ij}+(n-1)\delta},\,f^{\prime}_{\delta;\,\alpha_{i-1}}]\!],$ (5.13)
@@
$f_{(\delta-\alpha_{ij})+n\delta}=\frac{(-1)^{[\alpha_{i-1}]}}{[(\,\alpha_{ij}\,|\,\alpha_{i-1}\,)]_{q}}[\![f^{\prime}_{\delta;\,\alpha_{i-1}},\,f_{(\delta-\alpha_{ij})+(n-1)\delta}]\!],$ (5.15)
@@
$e^{\prime}_{n\delta;\,\alpha_{i}}=(-1)^{[\alpha_{i}]}[\![e_{\alpha_{i}+(n-1)\delta},\,e_{\delta-\alpha_{i}}]\!],\qquad f^{\prime}_{n\delta;\,\alpha_{i}}=(-1)^{[\alpha_{i}]}[\![f_{\alpha_{i}+(n-1)\delta},\,f_{\delta-\alpha_{i}}]\!].$ (5.16)
The prime in the notation for the root vectors corresponding to the imaginary roots $n\delta$ and $-n\delta$, $n\in\mathbb{Z}_{>0}$ is explained by the fact that to construct the expression for the universal $R$-matrix one uses another set of root vectors corresponding to these roots. They are introduced by the functional equations
@@
are defined as formal power series in $u$.

### v3

@@ 1. Introduction
Let us start with the necessary information on the affine Lie superalgebra ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$. More information can be found in the papers [50, 51].
We denote by $\mathcal{L}(\mathfrak{sl}_{M|N})$ the loop superalgebra of the Lie superalgebra $\mathfrak{sl}_{M|N}$, by ~$\widetilde{\mathcal{L}}(\mathfrak{sl}_{M|N})$ its standard central extension by a one-dimensional center $\mathbb{C}c$, and by ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$ the Lie superalgebra obtained from ~$\widetilde{\mathcal{L}}(\mathfrak{sl}_{M|N})$ by adding a natural derivation $d$. By definition, as a vector space
@@
and we use as the Cartan subsuperalgebra of ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$ the subsuperspace [^5] Here we use the natural embedding of $\mathfrak{sl}_{M|N}$ into ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$.
@@
We assume that $[c]=[d]=\overline{0}$. Further, we define
@@
where $d_{0}=1$. It is also convenient to denote
$\widetilde{\mathfrak{h}}=\mathfrak{h}\oplus\mathbb{C}\,c$
and identify $\mathfrak{h}$ and $\widetilde{\mathfrak{h}}$ with the corresponding subspaces of the subsuperalgebra $\widehat{\mathfrak{h}}$.
We identify the space $\mathfrak{h}^{*}$ with the subspace of $\widetilde{\mathfrak{h}}^{*}$ which consists of the elements $\lambda\in\widetilde{\mathfrak{h}}^{*}$ satisfying the equation
$\langle\lambda,\,c\rangle=0.$ (4.1)
Here and henceforth, we mark the elements of this subspace by a tilde. Explicitly, the identification is performed as follows. The element $\widetilde{\lambda}\in\widetilde{\mathfrak{h}}^{*}$ satisfying equation (4.1) is identified with the element $\lambda\in\mathfrak{h}^{*}$ defined by the equations
$\langle\lambda,\,h_{i}\rangle=\langle\widetilde{\lambda},\,h_{i}\rangle,\qquad i=1,\ldots,M+N-1.$
In the opposite direction, given an element $\lambda\in\mathfrak{h}^{*}$, identify it with the element $\widetilde{\lambda}\in\widetilde{\mathfrak{h}}^{*}$ determined by the relations
$\langle\widetilde{\lambda},\,h_{0}\rangle=-\ \sum_{i=1}^{\mathclap{M+N-1}}\ \langle\lambda,\,h_{i}\rangle,\qquad\langle\widetilde{\lambda},\,h_{i}\rangle=\langle\lambda,\,h_{i}\rangle,\quad i=1,\ldots,M+N-1.$
It is clear that the element $\widetilde{\lambda}$ defined in this way satisfies equation (4.1).
After all, denote by $\delta$ the element of the space $\widehat{\mathfrak{h}}^{*}$ determined by the equations
@@
The system of positive roots of the Lie superalgebra ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$ is
@@
Fix a non-degenerate symmetric bilinear form on $\widehat{\mathfrak{h}}$ by the equations
@@
The quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is a unital associative $\mathbb{C}$-superalgebra generated by the elements
$q^{x},\quad x\in\widetilde{\mathfrak{h}},\qquad e_{i},\quad f_{i},\quad i=0,1,\ldots,M+N-1.$
The $\mathbb{Z}_{2}$-grading of the superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is defined on generators as
@@
is called the root lattice of the Lie superalgebra $\mathfrak{gl}_{M|N}$. Assuming that
@@
we endow $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ with a $\widehat{Q}$-grading. We define the $q$-supercommutator for the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ in the same way as it was done for the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ in section 3.
The defining relations of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ look as [20]
@@
$[\![e_{i},\,e_{j}]\!]=0,\qquad[\![f_{j},\,f_{i}]\!]=0,\qquad\mbox{ if }(\,\alpha_{i}\,|\,\alpha_{j}\,)=0,$ (4.5)
$[\![e_{i},\,[\![e_{i},\,e_{i-1}]\!]]\!]=0,\qquad[\![[\![f_{i-1},\,f_{i}]\!],\,f_{i}]\!]=0,\qquad\mbox{ if }(\,\alpha_{i}\,|\,\alpha_{i}\,)\neq 0,$ (4.6)
$[\![e_{i},\,[\![e_{i},\,e_{i+1}]\!]]\!]=0,\qquad[\![[\![f_{i+1},\,f_{i}]\!],\,f_{i}]\!]=0,\qquad\mbox{ if }(\,\alpha_{i}\,|\,\alpha_{i}\,)\neq 0,$ (4.7)
$[\![[\![[\![e_{M-1},\,e_{M}]\!],\,e_{M+1}]\!],\,e_{M}]\!]=0,\qquad[\![f_{M},\,[\![f_{M+1},\,[\![f_{M},\,f_{M-1}]\!]]\!]]\!]=0,$ (4.8)
$[\![[\![[\![e_{1},\,e_{0}]\!],\,e_{M+N-1}]\!],\,e_{0}]\!]=0,\qquad[\![f_{0},\,[\![f_{M+N-1},\,[\![f_{0},\,f_{1}]\!]]\!]]\!]=0.$ (4.9)
The necessary periodicity with respect to the indices is implied in the relations (4.6) and (4.7). In the case of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{2|1}))$ instead of the relations (4.8) and (4.9) we have the relations
@@
$[\![[\![[\![[\![f_{2},\,f_{1}]\!],\,f_{0},]\!]\,f_{2}]\!],\,f_{0}]\!]=[\![[\![[\![[\![f_{0},\,f_{1}]\!],\,f_{2},]\!]\,f_{0}]\!],\,f_{2}]\!].$
From the point of view of quantum integrable systems, it is important that the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is a Hopf algebra with respect to an appropriately defined co-multiplication $\Delta$, antipode $S$ and counit $\varepsilon$. Note that the definition of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is invariant under the replacement of $q$ by $q^{-1}$, in contrast to the definition of the Hopf algebra structure. To eliminate the arising ambiguity, we give the explicit form of the Hopf algebra operations adopted in this paper,
@@ 5. Cartan–Weyl generators
Having in mind the $\widehat{Q}$-grading of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$, we say that an element $a\in\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$ is a root vector corresponding to a root $\gamma\in\widehat{\Delta}$, if $a\in\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))_{\gamma}$. In particular, the generators $e_{i}$ and $f_{i}$ are root vectors corresponding to the roots $\alpha_{i}$ and $-\alpha_{i}$.
One can construct linearly independent root vectors corresponding to all roots from $\widehat{\Delta}\subset\widehat{Q}$, see, for example, the papers [24, 25, 26]. They are called the Cartan–Weyl generators of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. If some ordering of roots is chosen, then the appropriately ordered monomials constructed from these vectors form a Poincaré–Birkhoff–Witt basis of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. To construct the root vectors we follow the procedure, based on a normal ordering of positive roots, proposed by Khoroshkin and Tolstoy [24, 25, 26].
@@
Assume that some normal ordering of positive roots is chosen. One says that a pair $(\alpha,\,\beta)$ of positive roots generates a root $\gamma$ if $\gamma=\alpha+\beta$ and $\alpha\prec\beta$. A pair of positive roots $(\alpha,\,\beta)$ generating a root $\gamma$ is said to be minimal if there is no other pair of positive roots $(\alpha^{\prime},\,\beta^{\prime})$ generating $\gamma$ such that $\alpha\prec\alpha^{\prime}\prec\beta^{\prime}\prec\alpha$.
It is convenient to denote a root vector corresponding to a positive root $\gamma$ by $e_{\gamma}$, and a root vector corresponding to a negative root $-\gamma$ by $f_{\gamma}$. Following the papers [24, 25, 26], we define root vectors by the following inductive procedure. Given a nonsimple positive root $\gamma\in\widehat{\Delta}_{+}$, let $(\alpha,\,\beta)$ be a minimal pair of positive roots generating $\gamma$. Then, if the root vectors $e_{\alpha}$, $e_{\beta}$ and $f_{\alpha}$, $f_{\beta}$ are already constructed, we define the root vectors $e_{\gamma}$ and $f_{\gamma}$ as
$e_{\gamma}\propto[\![e_{\alpha}\,,\,e_{\beta}]\!],\qquad f_{\gamma}\propto[\![f_{\beta}\,,\,f_{\alpha}]\!].$
@@
We use the normal order of the positive roots of the superalgebra Lie ^$\widehat{\mathcal{L}}(\mathfrak{sl}_{M|N})$ defined as follows, see also [52]. First we put $\alpha_{ij}+m\delta\prec\alpha_{ij}+n\delta$ if $m<n$, and $(\delta-\alpha_{ij})+m\delta\prec(\delta-\alpha_{ij})+n\delta$ if $m>n$. Then $\alpha_{ij}+m\delta\prec\alpha_{kl}+n\delta$ if $i<k$, or if $i=k$ and $j<l$. Further, $(\delta-\alpha_{ij})+m\delta\prec(\delta-\alpha_{kl})+n\delta$ if $i>k$, or, if $i=k$ and $j<l$. Finally, we assume that the relation (5.1) is valid.
@@
$e_{\alpha_{ij}}=[\![e_{\alpha_{i,\,j-1}},\,e_{\alpha_{j-1,\,j}}]\!],\qquad f_{\alpha_{ij}}=[\![f_{\alpha_{j-1,\,j}},\,f_{\alpha_{i,\,j-1}}]\!]$ (5.2)
for $j>i+1$. Further, taking into account that $\alpha_{0}=\delta-\alpha_{1,\,M+N}$, we put
@@
$e_{\delta-\alpha_{i,\,M+N}}=[\![e_{\alpha_{i-1,\,i}},\,e_{\delta-\alpha_{i-1,\,M+N}}]\!],\qquad f_{\delta-\alpha_{i,\,M+N}}=[\![f_{\delta-\alpha_{i-1,\,M+N}},\,f_{\alpha_{i-1,\,i}}]\!]$ (5.4)
@@
$e_{\delta-\alpha_{ij}}=[\![e_{\alpha_{j,\,j+1}},\,e_{\delta-\alpha_{i,\,j+1}}]\!],\qquad f_{\delta-\alpha_{ij}}=[\![f_{\delta-\alpha_{i,\,j+1}},\,f_{\alpha_{j,\,j+1}}]\!]$ (5.5)
for $j<M+N$. The root vectors corresponding to the roots $\delta$ and $-\delta$ are indexed by the simple roots $\alpha_{i}$, $i=1,\ldots,M+N-1$, and are defined by the relations
$e^{\prime}_{\delta;\,\alpha_{i}}=(-1)^{[\alpha_{i}]}[\![e_{\alpha_{i}},\,e_{\delta-\alpha_{i}}]\!],\qquad f^{\prime}_{\delta;\,\alpha_{i}}=(-1)^{[\alpha_{i}]}[\![f_{\delta-\alpha_{i}},\,f_{\alpha_{i}}]\!].$ (5.6)
@@
$f_{\alpha_{ij}+n\delta}=\frac{(-1)^{[\alpha_{i}]}}{[(\,\alpha_{ij}\,|\,\alpha_{i}\,)]_{q}}[\![f^{\prime}_{\delta;\,\alpha_{i}}\,f_{\alpha_{ij}+(n-1)\delta}]\!],$ (5.9)
@@
$f_{(\delta-\alpha_{ij})+n\delta}=\frac{(-1)^{[\alpha_{i}]}}{[(\,\alpha_{ij}\,|\,\alpha_{i}\,)]_{q}}[\![f_{(\delta-\alpha_{ij})+(n-1)\delta},\,f^{\prime}_{\delta;\,\alpha_{i}}]\!],$ (5.11)
@@
$f_{\alpha_{ij}+n\delta}=\frac{(-1)^{[\alpha_{i-1}]}}{[(\,\alpha_{ij}\,|\,\alpha_{i-1}\,)]_{q}}[\![f^{\prime}_{\delta;\,\alpha_{i-1}},f_{\alpha_{ij}+(n-1)\delta}]\!],$ (5.13)
@@
$f_{(\delta-\alpha_{ij})+n\delta}=\frac{(-1)^{[\alpha_{i-1}]}}{[(\,\alpha_{ij}\,|\,\alpha_{i-1}\,)]_{q}}[\![f_{(\delta-\alpha_{ij})+(n-1)\delta},\,f^{\prime}_{\delta;\,\alpha_{i-1}}]\!],$ (5.15)
@@
$e^{\prime}_{n\delta;\,\alpha_{i}}=(-1)^{[\alpha_{i}]}[\![e_{\alpha_{i}+(n-1)\delta},\,e_{\delta-\alpha_{i}}]\!],\qquad f^{\prime}_{n\delta;\,\alpha_{i}}=(-1)^{[\alpha_{i}]}[\![f_{\delta-\alpha_{i}},\,f_{\alpha_{i}+(n-1)\delta}]\!].$ (5.16)
The prime in the notation for the root vectors corresponding to the imaginary roots is explained by the fact that to construct the expression for the universal $R$-matrix one uses another set of root vectors corresponding to these roots. They are introduced by the functional equations
@@
are defined as formal power series in $u$ and $u^{-1}$, respectively.